FOURIER TRANSFORM EMISSION SPECTROSCOPY OF NEW ELECTRONIC TRANSITIONS OF VN AND VO

\underline{R. S. RAM}, Department of Chemistry, University of Arizona, Tucson, AZ 85721; P. F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario, Canada N2L 3G1; S. P. DAVIS, Department of Physics, University of California, Berkeley, California 94720.